OpenStudy (anonymous):

HELP! How would I solve this equation 2x+9=3-x

OpenStudy (anonymous):

x= -2

OpenStudy (anonymous):

She asked how. Please do not give direct answers! This should help: Let's solve your equation step-by-step. 2x+9=3−x Step 1: Simplify both sides of the equation. 2x+9=3−x Simplif 2x+9=−x+3 Step 2: Add x to both sides. 2x+9+x=−x+3+x 3x+9=3 Step 3: Subtract 9 from both sides. 3x+9−9=3−9 3x=−6 Step 4: Divide both sides by 3. 3x 3 = −6 3 x=−2 Answer: x=−2